55 square yards. The width is 7 yards, find the length
Answer by cleomenius(959)

I assume 55 is the perimeter.
7 + 7 = 14
55 - 14 = 41
41 / 2 = 20.5 yards, the length.
2(20.5) + 2 (7) = 55 yards.
Cleomenius.